# Diffusion Models Are Real-Time Game Engines

### v1

@@ Abstract
We present GameNGen, the first game engine powered entirely by a neural model that enables real-time interaction with a complex environment over long trajectories at high quality. GameNGen can interactively simulate the classic game DOOM at over 20 frames per second on a single TPU. Next frame prediction achieves a PSNR of 29.4, comparable to lossy JPEG compression. Human raters are only slightly better than random chance at distinguishing short clips of the game from clips of the simulation. GameNGen is trained in two phases: (1) an RL-agent learns to play the game and the training sessions are recorded, and (2) a diffusion model is trained to produce the next frame, conditioned on the sequence of past frames and actions. Conditioning augmentations enable stable auto-regressive generation over long trajectories.
See https://gamengen.github.io for demo videos.
@@ 1 Introduction
Computer games are manually crafted software systems centered around the following game loop: (1) gather user inputs, (2) update the game state, and (3) render it to screen pixels. This game loop, running at high frame rates, creates the illusion of an interactive virtual world for the player. Such game loops are classically run on standard computers, and while there have been many amazing attempts at running games on bespoke hardware (e.g. the iconic game DOOM has been run on kitchen appliances such as a toaster and a microwave, a treadmill, a camera, an iPod, and within the game of Minecraft, to name just a few examples111See https://www.reddit.com/r/itrunsdoom/), in all of these cases the hardware is still emulating the manually written game software as-is. Furthermore, while vastly different game engines exist, the game state updates and rendering logic in all are composed of a set of manual rules, programmed or configured by hand.
@@
In this work we demonstrate that the answer is yes. Specifically, we show that a complex video game, the iconic game DOOM, can be run on a neural network (an augmented version of the open Stable Diffusion v1.4 (Rombach et al., 2022)), in real-time, while achieving a visual quality comparable to that of the original game. While not an exact simulation, the neural model is able to perform complex game state updates, such as tallying health and ammo, attacking enemies, damaging objects, opening doors, and persist the game state over long trajectories.
GameNGen answers one of the important questions on the road towards a new paradigm for game engines, one where games are automatically generated, similarly to how images and videos are generated by neural models in recent years. Key questions remain, such as how these neural game engines would be trained and how games would be effectively created in the first place, including how to best leverage human inputs. We are nevertheless extremely excited for the possibilities of this new paradigm.
@@ 2 Interactive World Simulation
An Interactive Environment consists of a space of latent states , a space of partial projections of the latent space , a partial projection function , a set of actions , and a transition probability function such that .
For example, in the case of the game DOOM, is the program’s dynamic memory contents, is the rendered screen pixels, is the game’s rendering logic, is the set of key presses and mouse movements, and is the program’s logic given the player’s input (including any potential non-determinism).
@@ 3 GameNGen
GameNGen (pronounced “game engine”) is a generative diffusion model that learns to simulate the game under the settings of Section 2. In order to collect training data for this model, with the teacher forcing objective, we first train a separate model to interact with the environment. The two models (agent and generative) are trained in sequence. The entirety of the agent’s actions and observations corpus during training is maintained and becomes the training dataset for the generative model in a second stage. See Figure 3.
@@ 3.1 Data Collection via Agent Play
We record the agent’s training trajectories throughout the entire training process, which includes different skill levels of play. This set of recorded trajectories is our dataset, used for training the generative model (see Section 3.2).
@@ 3.2 Training the Generative Diffusion Model
We re-purpose a pre-trained text-to-image diffusion model, Stable Diffusion v1.4 (Rombach et al., 2022). We condition the model on trajectories , i.e. on a sequence of previous actions and observations (frames) and remove all text conditioning. Specifically, to condition on actions, we simply learn an embedding from each action (e.g. a specific key press) into a single token and replace the cross attention from the text into this encoded actions sequence. In order to condition on observations (i.e. previous frames) we encode them into latent space using the auto-encoder and concatenate them in the latent channels dimension to the noised latents (see Figure 3). We also experimented conditioning on these past observations via cross-attention but observed no meaningful improvements.
We train the model to minimize the diffusion loss with velocity parameterization (Salimans & Ho, 2022b):
@@ 3.2.1 Mitigating Auto-Regressive Drift Using Noise Augmentation
The domain shift between training with teacher-forcing and auto-regressive sampling leads to error accumulation and fast degradation in sample quality, as demonstrated in Figure 4. To avoid this divergence due to auto-regressive application of the model, we corrupt context frames by adding a varying amount of Gaussian noise to encoded frames in training time, while providing the noise level as input to the model, following Ho et al. (2021). To that effect, we sample a noise level uniformly up to a maximal value, discretize it and learn an embedding for each bucket (see Figure 3). This allows the network to correct information sampled in previous frames, and is critical for preserving frame quality over time. During inference, the added noise level can be controlled to maximize quality, although we find that even with no added noise the results are significantly improved. We ablate the impact of this method in section 5.2.2.
@@ 3.2.2 Latent Decoder Fine-tuning
The pre-trained auto-encoder of Stable Diffusion v1.4, which compresses 8x8 pixel patches into 4 latent channels, results in meaningful artifacts when predicting game frames, which affect small details and particularly the bottom bar HUD (“heads up display”). To leverage the pre-trained knowledge while improving image quality, we train just the decoder of the latent auto-encoder using an MSE loss computed against the target frame pixels. It might be possible to improve quality even further using a perceptual loss such as LPIPS (Zhang et al. (2018)), which we leave to future work. Importantly, note that this fine tuning process happens completely separately from the U-Net fine-tuning, and that notably the auto-regressive generation isn’t affected by it (we only condition auto-regressively on the latents, not the pixels). Appendix A.2 shows examples of generations with and without fine-tuning the auto-encoder.
@@ 3.3.1 Setup
We also experimented with generating 4 samples in parallel and combining the results, with the hope of preventing rare extreme predictions from being accepted and to reduce error accumulation. We experimented both with averaging the samples and with choosing the sample closest to the median. Averaging performed slightly worse than single frame, and choosing the closest to the median performed only negligibly better. Since both increase the hardware requirements to 4 TPUs, we opt to not use them, but note that this might be an interesting area for future work.
@@ 3.3.2 Denoiser Sampling Steps
During inference, we need to run both the U-Net denoiser (for a number of steps) and the auto-encoder. On our hardware configuration (a TPU-v5), a single denoiser step and an evaluation of the auto-encoder both takes 10ms. If we ran our model with a single denoiser step, the minimum total latency possible in our setup would be 20ms per frame, or 50 frames per second. Usually, generative diffusion models, such as Stable Diffusion, don’t produce high quality results with a single denoising step, and instead require dozens of sampling steps to generate a high quality image. Surprisingly, we found that we can robustly simulate DOOM, with only 4 DDIM sampling steps (Song et al., 2020). In fact, we observe no degradation in simulation quality when using 4 sampling steps vs 20 steps or more (see Appendix A.4).
@@
Since we do observe degradation when using just a single sampling step, we also experimented with model distillation similarly to (Yin et al., 2024; Wang et al., 2023) in the single-step setting. Distillation does help substantially there (allowing us to reach 50 FPS as above), but still comes at a some cost to simulation quality, so we opt to use the 4-step version without distillation for our method (see Appendix A.4). This is an interesting area for further research.
We note that it is trivial to further increase the image generation rate substantially by parallelizing the generation of several frames on additional hardware, similarly to NVidia’s classic SLI Alternate Frame Rendering (AFR) technique. Similarly to AFR, the actual simulation rate would not increase and input lag would not reduce.
@@ 4.1 Agent Training
The agent model is trained using PPO (Schulman et al., 2017), with a simple CNN as the feature network, following Mnih et al. (2015). It is trained on CPU using the Stable Baselines 3 infrastructure (Raffin et al., 2021). The agent is provided with downscaled versions of the frame images and in-game map, each at resolution 160x120. The agent also has access to the last 32 actions it performed. The feature network computes a representation of size 512 for each image. PPO’s actor and critic are 2-layer MLP heads on top of a concatenation of the outputs of the image feature network and the sequence of past actions. We train the agent to play the game using the Vizdoom environment (Wydmuch et al., 2019). We run 8 games in parallel, each with a replay buffer size of 512, a discount factor , and an entropy coefficient of . In each iteration, the network is trained using a batch size of 64 for 10 epochs, with a learning rate of 1e-4. We perform a total of 10M environment steps.
@@ 4.2 Generative Model Training
We train all simulation models from a pretrained checkpoint of Stable Diffusion 1.4, unfreezing all U-Net parameters. We use a batch size of 128 and a constant learning rate of 2e-5, with the Adafactor optimizer without weight decay (Shazeer & Stern, 2018) and gradient clipping of 1.0. We change the diffusion loss parameterization to be v-prediction (Salimans & Ho (2022a). The context frames condition is dropped with probability 0.1 to allow CFG during inference. We train using 128 TPU-v5e devices with data parallelization. Unless noted otherwise, all results in the paper are after 700,000 training steps. For noise augmentation (Section 3.2.1), we use a maximal noise level of 0.7, with 10 embedding buckets. We use a batch size of 2,048 for optimizing the latent decoder, other training parameters are identical to those of the denoiser. For training data, we use all trajectories played by the agent during RL training as well as evaluation data during training, unless mentioned otherwise. Overall we generate 900M frames for training. All image frames (during training, inference, and conditioning) are at a resolution of 320x240 padded to 320x256. We use a context length of 64 (i.e. the model is provided its own last 64 predictions as well as the last 64 actions).
@@ 5.1 Simulation Quality
Overall, our method achieves a simulation quality comparable to the original game over long trajectories in terms of image quality. For short trajectories, human raters are only slightly better than random chance at distinguishing between clips of the simulation and the actual game.
@@
Human Evaluation. As another measurement of simulation quality, we provided 10 human raters with 130 random short clips (of lengths 1.6 seconds and 3.2 seconds) of our simulation side by side with the real game. The raters were tasked with recognizing the real game (see Figure 14 in Appendix A.6). The raters only choose the actual game over the simulation in 58% or 60% of the time (for the 1.6 seconds and 3.2 seconds clips, respectively).
@@ 5.2.1 Context Length
We evaluate the impact of changing the number of past observations in the conditioning context by training models with (recall that our method uses ). This affects both the number of historical frames and actions. We train the models for 200,000 steps keeping the decoder frozen and evaluate on test-set trajectories from 5 levels. See the results in Table 1. As expected, we observe that generation quality improves with the length of the context. Interestingly, we observe that while the improvement is large at first (e.g. between 1 and 2 frames), we quickly approach an asymptote and further increasing the context size provides only small improvements in quality. This is somewhat surprising as even with our maximal context length, the model only has access to a little over 3 seconds of history. Notably, we observe that much of the game state is persisted for much longer periods (see Section 7). While the length of the conditioning context is an important limitation, Table 1 hints that we’d likely need to change the architecture of our model to efficiently support longer contexts, and employ better selection of the past frames to condition on, which we leave for future work.
@@ 6 Related Work
Interactive 3D Simulation
Simulating visual and physical processes of 2D and 3D environments and allowing interactive exploration of them is an extensively developed field in computer graphics (Akenine-Mller et al., 2018). Game Engines, such as Unreal and Unity, are software that processes representations of scene geometry and renders a stream of images in response to user interactions. The game engine is responsible for keeping track of all world state, e.g. the player position and movement, objects, character animation and lighting. It also tracks the game logic, e.g. points gained by accomplishing game objectives. Film and television productions use variants of ray-tracing (Shirley & Morley, 2008), which are too slow and compute-intensive for real time applications. In contrast, game engines must keep a very high frame rate (typically 30-60 FPS), and therefore rely on highly-optimized polygon rasterization, often accelerated by GPUs. Physical effects such as shadows, particles and lighting are often implemented using efficient heuristics rather than physically accurate simulation.
Neural 3D Simulation
Neural methods for reconstructing 3D representations have made significant advances over the last years. NeRFs (Mildenhall et al., 2020) parameterize radiance fields using a deep neural network that is specifically optimized for a given scene from a set of images taken from various camera poses. Once trained, novel point of views of the scene can be sampled using volume rendering methods. Gaussian Splatting (Kerbl et al., 2023) approaches build on NeRFs but represent scenes using 3D Gaussians and adapted rasterization methods, unlocking faster training and rendering times. While demonstrating impressive reconstruction results and real-time interactivity, these methods are often limited to static scenes.
Video Diffusion Models
Diffusion models achieved state-of-the-art results in text-to-image generation (Saharia et al., 2022; Rombach et al., 2022; Ramesh et al., 2022; Podell et al., 2023), a line of work that has also been applied for text-to-video generation tasks (Ho et al., 2022; Blattmann et al., 2023b; a; Gupta et al., 2023; Girdhar et al., 2023; Bar-Tal et al., 2024). Despite impressive advancement in realism, text adherence and temporal consistency, video diffusion models remain too slow for real-time applications. Our work extends this line of work and adapts it for real-time generation conditioned autoregressively on a history of past observations and actions.
Game Simulation and World Models
Several works attempted to train models for game simulation with actions inputs. Yang et al. (2023) build a diverse dataset of real-world and simulated videos and train a diffusion model to predict a continuation video given a previous video segment and a textual description of an action. Menapace et al. (2021) and Bruce et al. (2024) focus on unsupervised learning of actions from videos. Menapace et al. (2024) converts textual prompts to game states, which are later converted to a 3D representation using NeRF. Unlike these works, we focus on interactive playable real-time simulation, and demonstrate robustness over long-horizon trajectories. We leverage an RL agent to explore the game environment and create rollouts of observations and interactions for training our interactive game model. Another line of work explored learning a predictive model of the environment and using it for training an RL agent. Ha & Schmidhuber (2018) train a Variational Auto-Encoder (Kingma & Welling, 2014) to encode game frames into a latent vector, and then use an RNN to mimic the VizDoom game environment, training on random rollouts from a random policy (i.e. selecting an action at random). Then controller policy is learned by playing within the “hallucinated” environment. Hafner et al. (2020) demonstrate that an RL agent can be trained entirely on episodes generated by a learned world model in latent space. Also close to our work is Kim et al. (2020), that use an LSTM architecture for modeling the world state, coupled with a convolutional decoder for producing output frames and jointly trained under an adversarial objective. While this approach seems to produce reasonable results for simple games like PacMan, it struggles with simulating the complex environment of VizDoom and produces blurry samples. In contrast, GameNGen is able to generate samples comparable to those of the original game, see Figure 2. Finally, concurrently with our work, Alonso et al. (2024) train a diffusion world model to predict the next observation given observation history, and iteratively train the world model and an RL model on Atari games.
DOOM
@@
When DOOM released in 1993 it revolutionized the gaming industry. Introducing groundbreaking 3D graphics technology, it became a cornerstone of the first-person shooter genre, influencing countless other games. DOOM was studied by numerous research works. It provides an open-source implementation and a native resolution that is low enough for small sized models to simulate, while being complex enough to be a challenging test case. Finally, the authors have spent countless youth hours with the game. It was a trivial choice to use it in this work.

### v2

@@ Abstract
We present GameNGen, the first game engine powered entirely by a neural model that also enables real-time interaction with a complex environment over long trajectories at high quality. When trained on the classic game DOOM, GameNGen extracts gameplay and uses it to generate a playable environment that can interactively simulate new trajectories. GameNGen runs at 20 frames per second on a single TPU and remains stable over extended multi-minute play sessions. Next frame prediction achieves a PSNR of 29.4, comparable to lossy JPEG compression. Human raters are only slightly better than random chance at distinguishing short clips of the game from clips of the simulation, even after 5 minutes of auto-regressive generation. GameNGen is trained in two phases: (1) an RL-agent learns to play the game and the training sessions are recorded, and (2) a diffusion model is trained to produce the next frame, conditioned on the sequence of past frames and actions. Conditioning augmentations help ensure stable auto-regressive generation over long trajectories, and decoder fine-tuning improves the fidelity of visual details and text.
@@ 1 Introduction
Computer games are manually crafted software systems centered around the following game loop: (1) update the game state based on user input, and (2) render the game state to screen pixels. This game loop, running at high frame rates, creates the illusion of an interactive virtual world for the player. Such game loops are classically run on standard computers, and while there have been many impressive attempts at running games on bespoke hardware (e.g. the iconic game DOOM has been run on kitchen appliances, a treadmill, a camera, and within the game of Minecraft111See https://www.reddit.com/r/itrunsdoom/), in all of these cases the hardware is still emulating the manually written game software as-is. Furthermore, while vastly different game engines exist, the game state updates and rendering logic in all are composed of a set of manual rules, programmed or configured by hand.
@@
In this work we demonstrate that the answer is yes. Specifically, we show that a complex video game, the iconic game DOOM, can be run on a neural network, an augmented version of the open Stable Diffusion v1.4 (Rombach et al., 2022), in real-time, while achieving a visual quality comparable to that of the original game. While not an exact simulation (see limitations in Section 7), the neural model is able to perform complex game state updates, such as tallying health and ammo, attacking enemies, damaging objects, opening doors, and more generally persist the game state over long trajectories.
Our key contribution is a demonstration that a complex video game (DOOM) can be simulated by a neural network in real time with high quality on a single TPU. We provide concrete architecture and technical insights on how to (1) adapt a text-to-image diffusion model in a stable auto-regressive setup via noise augmentation, (2) achieve visual quality comparable to the original via fine-tuning the latent decoder, and (3) collect training data from an existing game at scale via an RL agent.
More broadly, demonstrating that real-time simulation of complex games on existing hardware is possible addresses one important question on the path towards a new paradigm for game engines – one where games are automatically generated, much like how images and videos have been generated by neural models in recent years. While bigger questions remain, such as how to use human input to create entirely new games instead of simulating existing ones, we are nevertheless excited for the possibilities of this new paradigm (see Section 7 for further discussion).
@@ 2 Interactive World Simulation
An Interactive Environment consists of a space of latent states , a space of observations of the latent space , a partial projection function , a set of actions , and a transition probability function such that .
For example, in the case of the game DOOM, is the program’s dynamic memory contents, is the rendered screen pixels, is the game’s rendering logic, is the set of key presses, and is the program’s logic given the player’s input (including any potential non-determinism).
@@ 3 GameNGen
GameNGen (pronounced “game engine”) is a generative diffusion model that learns to simulate the game under the settings of Section 2. In order to collect training data for this model at scale, with the teacher forcing objective, we first train a separate model to interact with the environment. The two models (agent and generative) are trained in sequence. The entirety of the agent’s actions and observations corpus during training is maintained and becomes the training dataset for the generative model in a second stage. See Figure 3.
@@ 3.1 Data Collection via Agent Play
We record the agent’s training trajectories throughout the entire training process, which includes different skill levels of play, starting with a random policy when the agent is untrained. This set of recorded trajectories is our dataset, used for training the generative model (see Section 3.2).
@@ 3.2 Training the Generative Diffusion Model
We re-purpose a pre-trained text-to-image diffusion model, Stable Diffusion v1.4 (Rombach et al., 2022) to predict the next frame in the game. We condition the model on trajectories , i.e. on a sequence of previous actions and observations (frames) and remove all text conditioning. Specifically, to condition on actions (i.e. key presses), we simply learn an embedding from each action into a single token and replace the cross attention from the text into this encoded actions sequence. In order to condition on observations (i.e. previous frames) we encode them into latent space using the auto-encoder and concatenate them in the latent channels dimension to the noised latents (see Figure 3). We also experimented with conditioning on these past observations via cross-attention but observed no meaningful improvements.
We train the model to minimize the diffusion loss with velocity parameterization (Salimans & Ho, 2022):
@@ 3.2.1 Mitigating Auto-Regressive Drift Using Noise Augmentation
The domain shift between training with teacher-forcing and auto-regressive sampling leads to error accumulation and fast degradation in sample quality, as demonstrated in Figure 4 (top). To avoid this divergence due to auto-regressive application of the model, we corrupt context frames by adding a varying amount of Gaussian noise to encoded frames in training time, while providing the noise level as input to the model, following Ho et al. (2021). To that effect, we sample a noise level uniformly up to a maximal value, discretize it and learn an embedding for each bucket (see Figure 3). This allows the network to correct information sampled in previous frames, and is critical for preserving frame quality over time. During inference, the added noise level can be controlled to maximize quality, although we find that even with no added noise the results are significantly improved. We ablate the impact of this method in Section 5.2.2.
@@ 3.2.2 Latent Decoder Fine-tuning
The pre-trained auto-encoder of Stable Diffusion v1.4, which compresses 8x8 pixel patches into 4 latent channels, results in meaningful artifacts when predicting game frames, which affect small details and particularly the bottom bar HUD (“heads up display”). To hopefully leverage some of the pre-trained knowledge while improving image quality, we train just the decoder of the latent auto-encoder using an MSE loss computed against the target frame pixels. Importantly, note that this fine tuning process happens completely separately from the U-Net fine-tuning, and that notably the auto-regressive generation isn’t affected by it (we only condition auto-regressively on the latents, not the pixels). Appendix A.2 shows examples of generations with and without fine-tuning the auto-encoder.
@@ 3.3.2 Denoiser Sampling Steps
During inference, we need to run both the U-Net denoiser (for a number of steps) and the auto-encoder. On our hardware configuration (a single TPU-v5), a single denoiser step and an evaluation of the auto-encoder both takes 10ms. If we ran our model with a single denoiser step, the minimum total latency possible in our setup would therefore be 20ms per frame, or 50 frames per second. Usually, generative diffusion models, such as Stable Diffusion, don’t produce high quality results with a single denoising step, and instead require dozens of sampling steps to generate a high quality image. Surprisingly, we found that we can robustly simulate DOOM, with only 4 DDIM sampling steps (Song et al., 2020). In fact, we observe no degradation in simulation quality when using 4 sampling steps vs 20 steps or more (see Table 1 and Appendix A.6).
| Steps | PSNR | LPIPS |
| D | ||
| 1 | ||
| 2 | ||
| 4 | ||
| 8 | ||
| 16 | ||
| 32 | ||
| 64 |
@@
Since we do observe degradation when using just a single sampling step, we also experimented with model distillation similarly to (Yin et al., 2024; Wang et al., 2023) in the single-step setting. Distillation does help substantially there (allowing us to reach 50 FPS as above), but still comes at a some cost to simulation quality, so we opt to use the 4-step version without distillation for our method (see Appendix A.6).
@@ 4.1 Agent Training
The agent model is trained using PPO (Schulman et al., 2017), with a simple CNN as the feature network, following Mnih et al. (2015b). It is trained on CPU using the Stable Baselines 3 infrastructure (Raffin et al., 2021). The agent is provided with downscaled versions of the frame images and in-game map, each at resolution 160x120. The agent also has access to the last 32 actions it performed. The feature network computes a representation of size 512 for each image. PPO’s actor and critic are 2-layer MLP heads on top of a concatenation of the outputs of the image feature network and the sequence of past actions. We train the agent to play the game using the ViZDoom environment (Wydmuch et al., 2019). We run 8 games in parallel, each with a replay buffer size of 512, a discount factor , and an entropy coefficient of . In each iteration, the network is trained using a batch size of 64 for 10 epochs, with a learning rate of 1e-4. We perform a total of 50M environment steps.
@@ 4.2 Generative Model Training
We train all simulation models from a pretrained checkpoint of Stable Diffusion 1.4, unfreezing all U-Net parameters. We use a batch size of 128 and a constant learning rate of 2e-5, with the Adafactor optimizer without weight decay (Shazeer & Stern, 2018) and gradient clipping of 1.0. The context frames condition is dropped with probability 0.1 to allow CFG during inference. We train using 128 TPU-v5e devices with data parallelization. Unless noted otherwise, all results in the paper are after 700,000 training steps. For noise augmentation (Section 3.2.1), we use a maximal noise level of 0.7, with 10 embedding buckets. We use a batch size of 2,048 for optimizing the latent decoder, other training parameters are identical to those of the denoiser. For training data, we use a random subset of 70M examples from the recorded trajectories played by the agent during RL training and evaluation (see Appendix A.3 for results with smaller datasets). All image frames (during training, inference, and conditioning) are at a resolution of 320x240 padded to 320x256. We use a context length of 64 (i.e. the model is provided its own last 64 predictions as well as the last 64 actions).
@@ 5.1 Simulation Quality
Overall, our method achieves a simulation quality comparable to the original game over long trajectories in terms of image quality. Human raters are only slightly better than random chance at distinguishing between short clips of the simulation and the actual game.
@@
Human Evaluation. As another measurement of simulation quality, we provided 10 human raters with 130 random short clips (of lengths 1.6 seconds and 3.2 seconds) of our simulation side by side with the real game. The raters were tasked with recognizing the real game (see Figure 17 in Appendix A.8). The raters only choose the actual game over the simulation in 58% or 60% of the time (for the 1.6 seconds and 3.2 seconds clips, respectively). To evaluate the impact of accumulated errors from auto-regressive generation, we generated 150 additional side-by-side comparisons using clips of 3 seconds in length after 5 to 10 minutes of gameplay. Despite the extended period of auto-regressive generation, raters still performed at chance level, identifying the real game only 50% of the time. While human raters struggled to distinguish between the simulation and real gameplay in short clips, it is important to note that the authors, who are familiar with the specific limitations of the simulation, could often identify the real game after just a few seconds of play. For qualitative assessment of longer multi-minute clips, refer to the videos in the supplementary material.
@@ 5.2.1 Context Length
We evaluate the impact of changing the number of past observations in the conditioning context by training models with (recall that our method uses ). This affects both the number of historical frames and actions. We train the models for 200,000 steps keeping the decoder frozen and evaluate on test-set trajectories from 5 levels. See the results in Table 2. As expected, we observe that generation quality improves with the length of the context. Interestingly, we observe that while the improvement is large at first (e.g. between 1 and 2 frames), we quickly approach an asymptote and further increasing the context size provides only small improvements in quality. This is somewhat surprising as even with our maximal context length, the model only has access to a little over 3 seconds of history. Notably, we observe that much of the game state is persisted for much longer periods (see Section 7). While the length of the conditioning context is an important limitation, Table 2 hints that we’d likely need to change the architecture or training scheme of our model to efficiently support longer contexts, and employ better selection of the past frames to condition on, which we leave for future work.
@@ 6 Related Work
Game Simulation and World Models. Several works attempted to train models for game simulation with actions inputs. Yang et al. (2023) build a diverse dataset of real-world and simulated videos and train a diffusion model to predict a continuation video given a previous video segment and a textual description of an action. Menapace et al. (2021) and Bruce et al. (2024) focus on unsupervised learning of actions from videos. Menapace et al. (2024) converts textual prompts to game states, which are later converted to a 3D representation using NeRF.
Another line of work explored learning a predictive model of the environment and using it for training an RL agent. Ha & Schmidhuber (2018) train a Variational Auto-Encoder (Kingma & Welling, 2014) to encode game frames into a latent vector, and then use an RNN to mimic the ViZDoom game environment, training on random rollouts from a random policy (i.e. selecting an action at random). Then controller policy is learned by playing within the “hallucinated” environment. Hafner et al. (2020) demonstrate that an RL agent can be trained entirely on episodes generated by a learned world model in latent space.
Also close to our work is Kim et al. (2020), that use an LSTM architecture for modeling the world state, coupled with a convolutional decoder for producing output frames and jointly trained under an adversarial objective.
Yan et al. (2021) apply a decoder-only transformer architecture to video generation using a discrete latent space learned by a VQ-VAE, and show that it can be used for action-conditioned video generation on the VizDoom simulator.
Hu et al. (2023) focuses on the problem of action-conditioned world modeling in the domain of driving. A multi-modal auto-regressive transformer acts as the world model and predicts the latent-space image token based on past image, text and action tokens. Then a diffusion video decoder translates the image tokens into a pixel-space video. The models are trained on large corpus of real-work driving data (420M unique images).
Finally, concurrently with our work, Alonso et al. (2024) train a diffusion world model to predict the next observation given observation history, and iteratively train the world model and an RL model on Atari games. More recent version of their work also included a high-res simulation of Counter-Strike, trained on 95 hours of human game play recording.
Auto-regressive Diffusion Models. Some recent work explored auto-regressive architectures of diffusion models. Chen et al. (2024) diverge from traditional diffusion architectures by allowing each token to have its own level of noise in each time step, reporting this approach to improve stability of video generation beyond the training horizon, when using a convolutional RNN backbone. Ruhe et al. (2024) also allow variable levels of noise for different tokens, using a sliding window denoising process. Such approaches are interesting to explore for real-time game simulation, and we leave this to future work.
@@
DOOM. When DOOM was released in 1993 it revolutionized the gaming industry. Introducing groundbreaking 3D graphics technology, it became a cornerstone of the first-person shooter genre, influencing countless other games. DOOM was studied by numerous research works. It provides an open-source implementation and a native resolution that is low enough for small sized models to simulate, while being complex enough to be a challenging test case. Finally, the authors have spent countless youth hours with the game. It was a trivial choice to use it in this work.